# Can Large Language Models Beat Wall Street? Unveiling the Potential of AI in Stock Selection

### v1

@@ 3.5 Signal Generation
The signal generation component, as the final stage in the MarketSenseAI pipeline (Figure 1), integrates insights from the news, fundamentals, price dynamics, and macroeconomic analysis components. This process results in a comprehensive investment recommendation for a specific stock, paired with a detailed rationale.
In essence, we argue that an investment decision for a stock ($I_{s}$) is a function of the most important developments for the underlying company available in some extent in the news ($N_{s}$), the company’s financial health ($F_{s}$), stock’s performance in relation to competitors and the market ($P_{s}$), as well as the broader macroeconomic environment (see Equation 1).
@@
We posit that the state-of-the-art LLM, GPT-4, possesses the requisite capacity to weigh and reason upon these different categories of data, as evidenced by its demonstrated proficiency in complex financial reasoning tasks (Callanan et al., 2023). In its operation, the GPT-4 model is prompted to adopt the role of an expert financial analyst. This approach employs a Chain of Thought methodology, guiding the model through a multi-step analysis to make a decision on portfolio allocation for the upcoming month (Wei et al., 2022). The prompt structure is as follows:
@@
The model’s output, structured in a concise format, includes a decision (”buy,” ”sell,” or ”hold”) along with a clear, step-by-step explanation of the reasoning behind this choice. The terms ”buy” and ”sell” are defined within the context of portfolio positioning (long and short positions, respectively), while ”hold” indicates no inclusion in the portfolio’s composition regarding the specific stock.
@@
HOLD
The analysis […] presents a mixed picture. Starting with the news analysis, while Apple remains a dominant player in the tech sector with solid financials and a successful product launch in South Korea, it faces challenges such as a sales slowdown, competitive pressures, and regulatory scrutiny. The Chinese government ban and sluggish sales in China and Japan are concerning, but analysts see the recent stock dip as a buying opportunity, suggesting underlying strength. The company’s strategic moves, such as potentially ending its credit card partnership with Goldman Sachs, indicate a focus on its services business, which is a positive sign. However, the legal challenges and geopolitical risks cannot be ignored. In terms of price dynamics, Apple has outperformed the S&P 500 […] strong risk-adjusted performance. However, the maximum drawdown and the high correlation with the market suggest that Apple is not immune to market downturns. The macro environment analysis reveals a disinflation trend in the US and potential rate cuts in 2024, which could be favorable for equities. However, the global economy is expected to grow below trend, and there is a neutral stance on developed market equities, suggesting caution. The positive outlook on quality fixed income and US stocks is encouraging, but the vulnerability of the US dollar and downside risks to employment add uncertainty. Fundamentals analysis shows strong profitability and revenue growth for Apple, with increased net income and gross profit. The company’s debt levels have risen, which warrants monitoring, but the positive cash flow generation and growth in assets and equity are reassuring. Considering the mixed signals […] a ”hold” decision is prudent[…]
@@
News 0.923573 0.016005 0.865069
0.967842
Price Dynamics 0.907652 0.013444 0.854822
0.942414
Fundamentals 0.849269 0.033589 0.769349
0.934139
Macro 0.802891 0.017896 0.752009
0.870194
@@ 4.1 Data
The assessment period spans from December 1, 2022, to November 30, 2023. During this time, MarketSenseAI utilized diverse datasets for its in-context learning processes:
@@
    News: A total of 136,136 articles published from December 1, 2022, to November 30, 2023, averaging 3.72 articles per day per stock, with a standard deviation of 5.69. The mean number of tokens per article was 817, with a standard deviation of 1161. These data produced 36,904 daily, company-specific, news summaries and 1,200 monthly progressive summaries.
@@
    Fundamentals: Financial data were gathered from 612 quarterly reports of the S&P 100 stocks, starting from the second quarter of 2022. This data set produced 511 unique fundamentals summaries, averaging about 5 per stock.
@@
    Descriptions: : Concise descriptions of each stock and its sector, employed by Algorithm 1 to identify similar stocks.
@@
    Prices: Historical daily stock prices (adjusted close) from January 1, 2022, to November 30, 2023, were analyzed to compute stock price dynamics. This data was employed by the Stock Price Dynamics component to produce one summary per month for each stock, totaling 1,200 summaries.
@@
    Macro: 132 investment reports (20-30 pages each) from major financial institutions published between April and November 2023, were analyzed by MarketDigest. For predictions made from January to March 2023, macroeconomic summaries were not available for the signal generation component, resulting in 8 macroeconomic summaries used for signal generation.
To evaluate the system, the ”Signal Generation” component was fed at the end of each month with the latest available summaries (news, fundamentals, price dynamics, macro). While macroeconomic summaries were identical for all stocks, fundamental summaries were updated only when a new quarterly report for a stock became available. News and price summaries, being more dynamic, provided updated stock-specific insights each month. The investment signals generated by MarketSenseAI were subsequently assessed based on the actual stock performance in the month following the signal. Out of the 1,200 signals produced (12 months x 100 stocks), a breakdown of these signals reveals 220 ”buy”, 972 ”hold”, and 8 ”sell” signals. At the time of writing (December, 2023), 1,100 of these signals have been evaluated. This limitation is due to the time frame of the study, which allows for the analysis of only those signals where subsequent market data is available to measure their accuracy.
@@ 4.2.2 Market Performance
Table 8: Evaluated Investments Strategies on S&P 100 Stocks
Abbr. Description
MS
Equally weighted portfolio rebalanced monthly based on both ”buy” and ”sell” signals of MarketSenseAI.
MS-L
Equally weighted portfolio rebalanced monthly based on the ”buy” signals of MarketSenseAI.
MS-L-Cap
Capitalization-weighted portfolio rebalanced monthly based on the ”buy” signals of MarketSenseAI.
MS-Top10-SR
Equally weighted portfolio rebalanced monthly based on the 10 stocks with the best Sharpe Ratio of all the stocks with a ”buy” signal.
S&P100-Eq
Equally weighted portfolio of all the stocks of the S&P 100 index.
S&P100
Capitalization-weighted S&P 100 index (OEX ETF).
Naive
Equally weighted portfolio rebalanced monthly for all S&P 100 stocks with price above their corresponding 200 day moving average, fully allocated.
Naive-Top10
Equally weighted portfolio rebalanced monthly based on the 10 stocks with the best Sharpe Ratio and their price above their corresponding 200 day moving average, fully allocated.
MS-TopN-GPT
Equally weighted portfolio rebalanced monthly based on the $N$ stocks with the best score produced by GPT-4 after processing of all the stocks with a ”buy” signal.
MS-High-GPT
Equally weighted portfolio rebalanced monthly based on stocks with score grater than 7.5/10 produced by GPT-4 after processing of all the stocks with a ”buy” signal.
MS-Low-GPT
Equally weighted portfolio rebalanced monthly based on stocks with score lower or equal than 7.5/10 produced by GPT-4 after processing of all the stocks with a ”buy” signal.
@@
Total Return
The portfolio cumulative returns over a specific period (Equation 2)
Sharpe Ratio
A measure of risk-adjusted return; calculated as the average return earned in excess of the risk-free rate per unit of volatility.
Sortino Ratio
Similar to the Sharpe Ratio, but measures returns relative to downside risk, focusing on negative asset volatility.
Volatility
A statistical measure of the dispersion of returns for a given security or market index, measured using standard deviation.
Win Rate
The percentage of trades that are profitable out of the total number executed.
Maximum Drawdown
The maximum observed loss from a peak to a trough of a portfolio, before a new peak is attained.
Maximum Drawdown Duration
The length of time (days) for which the maximum drawdown is observed, measured from peak to recovery.
@@ 4.3 Setup
The implementation of MarketSenseAI was executed using Python 3.11, leveraging the LangChain framework (Chase, 2022) for prompt construction and utilizing OpenAI’s API for accessing the GPT-4 model. Each component of MarketSenseAI, as outlined in Section 3, functions independently, running as a standalone script. The outputs from these components are systematically stored in a datastore, ensuring organized and efficient data management.
@@ 5.1 Bootstrapping Evaluation Results
Detrend-Buy 6.04 95.99 51.26 96.28
Buy 21.20 97.39 53.77 74.47
Detrend-Sell 24.82 100 100 100
Sell 14.98 100 83.33 100
Detrend-Both 7.55 99.36 52.68 92.69
Both 21.86 100 54.63 99.45
@@
A particularly key observation is the marked increase in the hit ratio quantile for the ”Buy” signals following detrending. Considering the upward trend of the market during the evaluation period, this indicates that MarketSenseAI’s recommendations have a greater probability of success compared to a random signal generation approach. This finding is significant as it highlights the model’s capability in effectively pinpointing potential market-outperforming opportunities.
@@ 5.2 Market Performance Evaluation Results
It is important to note that the year 2023 witnessed disparate performances among stocks. While technology giants and AI-centric companies experienced a robust year, others showed modest returns (Thompson, 2023). Our analysis places special emphasis on equal-weighted indices, which helps counterbalance this disparity and illustrates MarketSenseAI’s potential. Additionally all results presented below account for transaction costs to assess real-world applicability and effectiveness of MarketSenseAI-derived strategies.
@@ 5.2.1 Vanilla strategies
The assessment of MarketSenseAI’s vanilla strategies, as detailed in Table 11 and illustrated in Figure 9, reveals the efficacy of LLM-driven investment strategies. The strategy following the complete set of MarketSenseAI’s signals (MS) equally weighted, yields a total return of 22.72% with a Sharpe and Sortino ratio of 2.02 and 3.13, respectively. The long-only version (MS-L) that takes into account only the ”buy” signals of MarketSenseAI gives similar results (somewhat lower) given the relatively few ’sell’ signals generated.
These results significantly outperform the equally-weighted S&P 100 (S&P100-Eq) both in total and risk-adjusted returns, with an ”alpha” of approximately 9%. The naive trend-following strategy (Naive), often used by market participants, yielded significantly lower results in 2023.
In terms of capitalization-weighted performance, the MS-L-Cap emerges as a top performer in terms of Sharpe and Sortino ratios, as well as total return, reaching an impressive 40.27% total return. This is a significant outperformance of 12.6% to the S&P100 ETF as depicted in Figure 10.
@@
Strategy Total Return(%) Sharpe Sortino Vol(%) Total Trades Win Rate(%) Max Ddn(%) Max Ddn Dur.
MS 22.72 (21.34) 2.02 3.13 16.74 160 63.31 7.94 96
MS-L 21.20 (20.52) 1.85 2.86 17.22 159 60.87 9.00 96
S&P100-Eq 13.57 (13.48) 1.38 2.09 15.37 622 72.61 10.66 90
Naive 5.88 (5.20) 0.66 0.94 15.45 469 58.50 11.00 103
MS-L-Cap 40.27 (39.56) 2.53 4.19 22.09 158 59.12 9.36 82
S&P100 27.71 (27.65) 2.39 3.79 16.73 1 N/A 9.24 77
@@
Strategy Total Return(%) Sharpe Sortino Vol(%) Total Trades Win Rate(%) Max Ddn(%) Max Ddn Dur.
MS-Top10-SR 13.33 (12.63) 1.08 1.57 20.29 101 64.84 12.66 65
MS-Top5-GPT 35.99 (35.09) 2.17 3.33 23.70 47 66.67 9.24 83
MS-Top10-GPT 30.91 (30.07) 2.28 3.57 19.49 100 67.78 10.77 58
MS-High-GPT 37.19 (36.11) 2.27 3.70 23.20 78 57.75 10.82 95
MS-Low-GPT 11.91 (10.95) 1.14 1.72 16.87 107 59.14 9.05 91
Naive-Top10 20.47 (19.88) 1.46 2.16 21.80 98 65.91 9.79 78
@@
5.2.2 Rank-based strategies
The strategies incorporating GPT-4 rankings significantly outperform other variants (MS-Top10-SR, Naive-Top10) with very strong total returns and risk adjusted metrics. Another interesting result is the significant outperformance of MS-High-GPT vs MS-Low-GPT strategy that underscores the value of the explanations attached to MarketSenseAI’s signals and a promising alternative to conventional risk/return-based stock ranking methods.

### v2

@@ 3.5 Signal Generation
The signal generation component, as the final stage in the MarketSenseAI pipeline (Figure 1), integrates the textual outputs from the news, fundamentals, price dynamics, and macroeconomic analysis components. This process results in a comprehensive investment recommendation for a specific stock, paired with a detailed rationale.
In essence, we argue that an investment decision for a stock is a function of the most important developments for the underlying company available in some extent in the news, the company’s financial health, stock’s performance in relation to competitors and the market, as well as the broader macroeconomic environment. The decision model is represented as:
@@
Where, $N_{s},F_{s},P_{s}$, and $M$ are stock-specific ($s$) textual representations of current news, fundamentals, price dynamics, and macroeconomic conditions available from the components present in sections 3.1-3.4.
We posit that the state-of-the-art LLM, GPT-4, possesses the requisite capacity to weigh and reason upon these different categories of data, as evidenced by its demonstrated proficiency in complex financial reasoning tasks (Callanan et al., 2023). In its operation, the GPT-4 model is prompted to adopt the role of an expert financial analyst. This approach employs a Chain of Thought methodology (Wei et al., 2022), guiding the model through a logical, multi-step reasoning process that reflects an expert financial analyst’s thinking pattern. By applying this technique, MarketSenseAI can effectively analyze and synthesize news, company fundamentals, stock performance data, and macroeconomic factors that could influence the given stock, thus providing reasoned and structured insights into stock selection. This approach is particularly useful in complex domains like finance, where the ability to navigate through multifaceted data and reason like an expert is crucial. Concurrently, in-context learning is employed to dynamically adjust the analysis based on current financial situations and evolving market data (Dong et al., 2022). This dual strategy allows MarketSenseAI to provide deep insights that adapt in changing market conditions and investors’ preferences, representing a significant advancement in AI-driven financial analysis. The prompt structure is as follows:
@@
The model’s output, structured in a concise format, includes a decision (”buy”, ”sell”, or ”hold”) along with a clear, step-by-step explanation of the reasoning behind this choice. The terms ”buy” and ”sell” are defined within the context of portfolio positioning (long and short positions, respectively), while ”hold” indicates no inclusion in the portfolio’s composition regarding the specific stock.
@@
HOLD The analysis […] presents a mixed picture. Starting with the news analysis, while Apple remains a dominant player in the tech sector with solid financials and a successful product launch in South Korea, it faces challenges such as a sales slowdown, competitive pressures, and regulatory scrutiny. The Chinese government ban and sluggish sales in China and Japan are concerning, but analysts see the recent stock dip as a buying opportunity, suggesting underlying strength. The company’s strategic moves, such as potentially ending its credit card partnership with Goldman Sachs, indicate a focus on its services business, which is a positive sign. However, the legal challenges and geopolitical risks cannot be ignored. In terms of price dynamics, Apple has outperformed the S&P 500 […] strong risk-adjusted performance. However, the maximum drawdown and the high correlation with the market suggest that Apple is not immune to market downturns. The macro environment analysis reveals a disinflation trend in the US and potential rate cuts in 2024, which could be favorable for equities. However, the global economy is expected to grow below trend, and there is a neutral stance on developed market equities, suggesting caution. The positive outlook on quality fixed income and US stocks is encouraging, but the vulnerability of the US dollar and downside risks to employment add uncertainty. Fundamentals analysis shows strong profitability and revenue growth for Apple, with increased net income and gross profit. The company’s debt levels have risen, which warrants monitoring, but the positive cash flow generation and growth in assets and equity are reassuring. Considering the mixed signals […] a ”hold” decision is prudent[…]
@@
News 0.923573 0.016005 0.865069 0.967842
Price Dynamics 0.907652 0.013444 0.854822 0.942414
Fundamentals 0.849269 0.033589 0.769349 0.934139
Macro 0.802891 0.017896 0.752009 0.870194
@@ 4.1 Data
The assessment period spans from December 1, 2022, to March 31, 2024. During this time, MarketSenseAI utilized diverse datasets for its in-context learning processes:
@@
    News: A total of 163,483 articles published from December 1, 2022, to February 29, 2024, averaging 4.57 articles per day per stock, with a standard deviation of 5.49. The mean number of tokens per article was 867, with a standard deviation of 1196. These data produced 35,229 daily, company-specific, news summaries and 1,500 monthly progressive summaries.
@@
    Fundamentals: Financial data were gathered from 612 quarterly reports of the S&P 100 stocks, starting from the second quarter of 2022. This data set produced 608 unique fundamentals summaries, averaging about 6 per stock.
@@
    Descriptions: Concise descriptions of each stock and its sector, employed by Algorithm 1 to identify similar stocks.
@@
    Prices: Historical daily stock prices (adjusted close) from January 1, 2022, to February 29, 2024, were analyzed to compute stock price dynamics. This data was employed by the Stock Price Dynamics component to produce one summary per month for each stock, totaling 1,500 summaries.
@@
    Macro: 187 investment reports (20-30 pages each) from major financial institutions published between April 2023 and February 2023, were analyzed by MarketDigest. For predictions made from January to March 2023, macroeconomic summaries were not available for the signal generation component, resulting in 11 macroeconomic summaries used for signal generation.
To evaluate the system, the ”Signal Generation” component was fed at the end of each month with the latest available summaries (news, fundamentals, price dynamics, macro). While macroeconomic summaries were identical for all stocks, fundamental summaries were updated only when a new quarterly report for a stock became available. News and price summaries, being more dynamic, provided updated stock-specific insights each month. The investment signals generated by MarketSenseAI were subsequently assessed based on the actual stock performance in the month following the signal. Overall, 1,500 signals produced (15 months x 100 stocks), a breakdown of these signals reveals 338 ”buy”, 1150 ”hold”, and 12 ”sell” signals.
@@ 4.2.2 Market Performance
Table 8: Evaluated Investments Strategies on S&P 100 Stocks
Abbreviation Description
MS Equally weighted portfolio rebalanced monthly based on both ”buy” and ”sell” signals of MarketSenseAI.
MS-L Equally weighted portfolio rebalanced monthly based on the ”buy” signals of MarketSenseAI.
MS-L-Cap Capitalization-weighted portfolio rebalanced monthly based on the ”buy” signals of MarketSenseAI.
MS-Top10-SR Equally weighted portfolio rebalanced monthly based on the 10 stocks with the best Sharpe Ratio of all the stocks with a ”buy” signal.
S&P100-Eq Equally weighted portfolio of all the stocks of the S&P 100 index.
S&P100 Capitalization-weighted S&P 100 index (OEF ETF).
Naive Equally weighted portfolio rebalanced monthly for all S&P 100 stocks with price above their corresponding 200 day moving average, fully allocated.
Naive-Top10 Equally weighted portfolio rebalanced monthly based on the 10 stocks with the best Sharpe Ratio and their price above their corresponding 200 day moving average, fully allocated.
MS-TopN-GPT Equally weighted portfolio rebalanced monthly based on the $N$ stocks with the best score produced by GPT-4 after processing of all the stocks with a ”buy” signal.
MS-High-GPT Equally weighted portfolio rebalanced monthly based on stocks with score grater than 7/10 produced by GPT-4 after processing of all the stocks with a ”buy” signal.
MS-Low-GPT Equally weighted portfolio rebalanced monthly based on stocks with score lower or equal than 7/10 produced by GPT-4 after processing of all the stocks with a ”buy” signal.
MS-TopN-Cap-GPT Capitalization-weighted portfolio rebalanced monthly based on the $N$ stocks with the best score produced by GPT-4 after processing of all the stocks with a ”buy” signal.
@@
Total Return The portfolio cumulative returns (%) over a specific period (Equation 8)
Sharpe Ratio A measure of risk-adjusted return; calculated as the average return earned in excess of the risk-free rate per unit of volatility.
Sortino Ratio Similar to the Sharpe Ratio, but measures returns relative to downside risk, focusing on negative asset volatility.
Volatility A statistical measure of the dispersion of returns for a given security or market index, measured using standard deviation.
Win Rate The percentage of trades that are profitable out of the total number executed.
Maximum Drawdown (Ddn) The maximum observed percentage loss from a peak to a trough of a portfolio, before a new peak is attained.
@@ 4.3 Setup
The implementation of MarketSenseAI was executed using Python 3.11, leveraging the LangChain framework (Chase, 2022) for prompt construction and utilizing OpenAI’s API for accessing the GPT-4 model. Each component of MarketSenseAI, as outlined in Section 3, functions independently, running as a standalone script. The outputs from these components are systematically stored in a datastore, ensuring organized and efficient data management. As MarketSenseAI accesses GPT-4 through OpenAI’s API, the operational aspects fall under OpenAI’s operational jurisdiction. This arrangement allows our framework to leverage the capabilities of GPT-4 efficiently, ensuring consistent processing times per stock, while offloading the computational and hardware management responsibilities to OpenAI’s robust infrastructure.
@@
To identify the most similar stocks of a given stock, a prerequisite for the Stock Price Dynamics Summarizer, we use the mpnet-base-v2 model available from Hugging Face’s Transformers framework and cosine_similarity function from Scikit-Learn Python library.
Experiments are run on a desktop computer with an AMD Ryzen 5 5600x 6-Core CPU, 32GiB of RAM, and an NVIDIA GeForce RTX 3070 GPU. For data storage a dedicated cloud based S3 bucket was used hosted on Amazon Web Services (AWS).
@@ 5.1 Bootstrapping Evaluation Results
Detrend-Buy 7.67 97.85 51.86 99.50
Buy 35.48 98.70 60.34 99.35
Detrend-Sell 22.62 100 72.73 100
Sell 0.10 100 63.64 100
Detrend-Both 8.41 99.55 52.61 94.20
Both 33.87 100 60.46 100
@@
A particularly key observation is the high hit ratio quantile for the ”Buy” signals following detrending. Considering the upward trend of the market during the evaluation period, this indicates that MarketSenseAI’s recommendations have a greater probability of success compared to a random signal generation approach. This finding is significant as it highlights the model’s capability in effectively pinpointing potential market-outperforming opportunities.
@@ 5.2 Market Performance Evaluation Results
It is important to note that the period of the experiments is characterized by disparate performances between the stocks. While technology giants and AI-centric companies experienced a robust year, others showed modest returns (Thompson, 2023). Our analysis places special emphasis on equal-weighted indices, which helps counterbalance this disparity and illustrates MarketSenseAI’s potential. Additionally all results presented below account for transaction costs to assess real-world applicability and effectiveness of MarketSenseAI-derived strategies.
@@ 5.2.1 Vanilla strategies
The assessment of MarketSenseAI’s vanilla strategies, as detailed in Table 11 and illustrated in Figure 9, reveals the efficacy of LLM-driven investment strategies. The strategy following the complete set of MarketSenseAI’s signals (MS) equally weighted, yields an impressive total return of 35.48% (32.94% after transaction costs) with a Sharpe and Sortino ratio of 2.49 and 3.87, respectively. The long-only version (MS-L) that takes into account only the ”buy” signals of MarketSenseAI gives similar results given the relatively few ”sell” signals generated.
These results significantly outperform the equally-weighted S&P 100 (S&P100-Eq) both in total and risk-adjusted returns, with an ”alpha” of approximately 10% and 28% higher Sortino ratio. The naive trend-following strategy (Naive), often used by market participants, yielded significantly lower results.
In terms of capitalization-weighted performance, the MS-L-Cap emerges as a top performer in terms of Sharpe and Sortino ratios, as well as total return, reaching an impressive 66% total return. This is a significant outperformance of 43% to the S&P100 ETF as depicted in Figure 10.
@@
Strategy Total Return(%) Sharpe Sortino Vol(%) Win Rate(%) Max Ddn(%)
MS 35.48 (32.94) 2.49 3.87 15.76 65.68 8.47
MS-L 35.79 (34.82) 2.41 3.75 16.44 65.02 9.00
S&P100-Eq 25.22 (25.12) 1.98 3.02 14.72 76.36 10.66
Naive 17.89 (17.45) 1.47 2.14 14.71 65.09 11.00
MS-L-Cap 66.22 (65.25) 2.90 4.95 22.81 65.88 9.64
S&P100 43.27 (43.20) 2.86 4.61 16.17 N/A 9.24
@@ 5.2.2 Rank-based strategies
Strategy Total Return(%) Sharpe Sortino Vol(%) Win Rate(%) Max Ddn(%)
MS-Top10-SR 23.13 (22.12) 1.45 2.11 19.27 67.8 12.66
MS-Top5-GPT 50.96 (49.67) 2.26 3.69 24.01 68.42 11.39
MS-Top10-GPT 49.09 (48.07) 2.68 4.29 19.37 74.1 7.66
MS-High-GPT 39.47 (38.35) 2.28 3.44 19.08 71.9 9.73
MS-Low-GPT 25.66 (24.27) 1.76 2.64 17.04 55.1 12.22
Naive-Top10 29.01 (28.18) 1.67 2.48 20.29 69.3 9.79
MS-Top10-Cap-GPT 72.87 (71.64) 2.80 4.89 25.61 71.2 10.77
@@
The MS-Top10-SR strategy, which selects the ”buy” stocks with the highest Sharpe ratio, delivered a total return of 23.13% (22.12% after transaction costs). However, strategies utilizing GPT-4 to rank stocks exhibit significantly superior performance, both in total and risk-adjusted returns, highlighting the advanced analytical capabilities of LLM in selecting high-potential investments as well as the importance of the information generated by MarketSenseAI.
The MS-Top5-GPT and MS-Top10-GPT strategies, which select the top 5 and top 10 GPT-4 ranked stocks respectively, achieved remarkable total returns of 50.96% (49.67% after costs) and 49.09% (48.07% after costs). These strategies not only surpassed the returns of MS-Top10-SR and the Naive-Top10 benchmarks but also demonstrated impressive Sharpe and Sortino ratios, indicating high returns per unit of risk.
Notably, the MS-Top10-GPT strategy also achieved the highest win rate of 74.1% and the lowest maximum drawdown of 7.66, further evidencing the robustness of GPT-4’s ranking mechanism in navigating market volatilities and capitalizing on growth opportunities.
The differentiation between MS-High-GPT and MS-Low-GPT strategies, yielding 39.47% (38.35% after costs) and 25.66% (24.27% after costs) respectively, elucidates the significant advantage of selecting stocks based on MarketSenseAI’s explanations. This distinction highlights the added value of AI ranking over conventional financial engineering methods that may focus solely on historical returns or volatility.
Furthermore, the MS-Top10-Cap-GPT strategy, which applies a capitalization-weighted approach to the top GPT-4 ranked stocks, stands out with a staggering total return of 72.87% (71.64% after costs), the highest among the strategies evaluated.